# Greedy Perspectives: Multi-Drone View Planning for Collaborative Perception in Cluttered Environments

### v2

@@ VII Conclusion and Future Work
In this work, we presented a novel system for multi-robot view planning based on sequential greedy planning with an occlusion-aware objective. Through evaluation in five different scenarios, we observe sequential planning outperforming formation-based planning and specifically excels in obstacle-dense environments. Additionally, we observe similar perception performance for sequential planning with and without inter-robot collision constraints. This demonstrates that sequential planning is able to find good solutions when accounting for possible collisions between robots (despite no longer satisfying requirements for bounded suboptimality). In future work, we aim to extend deploy this view planner to a 3D human pose reconstruction task, and optimize our implementation to run at real-time rates.

### v3

@@ VII Conclusion and Future Work
In this work, we presented a novel system for multi-robot view planning based on sequential greedy planning with an occlusion-aware objective. Through evaluation in five different scenarios, we observe sequential planning outperforming formation-based planning and specifically excelling in obstacle-dense environments. Additionally, we observe similar perception performance for sequential planning with and without inter-robot collision constraints. This demonstrates that sequential planning is able to find good solutions when accounting for possible collisions between robots (despite no longer satisfying requirements for bounded suboptimality). In future work, we aim to extend this view planner to a 3D human pose reconstruction task, and optimize our implementation to run at real-time rates.